OpenStudy (anonymous):

Josh works at $10.00 per hour. If he worked 45 hours this week, what is his gross pay if his job earns time and a half for overtime? $400.00 $450.00 $475.00 $500.00

OpenStudy (anonymous):

is this one 475.00 @BAdhi

OpenStudy (zpupster):

overtime starts after 40 hours 40*10.00=400.00 time and a half 10.00*1.5 = 15.00 5* 15.00 = 75.00 400 + 75 = 475